Prev Next

SW&IT Meeting Minutes 2019-12-16

Attendees: Carl, David, Igor, Matthieu, Petr, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20191118
LR20191118 - Bec, Matthieu 64.0
LR20191118 - Holmberg, Carl 64.0
LR20191118 - Colson, Andrew 0.0
LR20191118 - Edmeades, David 64.0
LR20191118 - Sola, Igor 64.0
LR20191118 - Kubanek, Petr 24.0
LR20191118 - Hooper, Stephen 64.0
LR20191118 - Zhang, Yang 61.0
LR20191118 - Zhang, Xianyu 64.0

20191202
LR20191202 - Bec, Matthieu 80.0
LR20191202 - Holmberg, Carl 80.0
LR20191202 - Colson, Andrew 40.0
LR20191202 - Edmeades, David 40.0
LR20191202 - Sola, Igor 45.5
LR20191202 - Kubanek, Petr 78.0
LR20191202 - Hooper, Stephen 64.0
LR20191202 - Zhang, Yang 80.0
LR20191202 - Zhang, Xianyu 72.0

News from management

  • Personnel announcement
  • LBTI/PEPSI/SHARK-VIS data into SciArc: not immediately

Action Items

Owner Action Due Date Completion Date

Open topics

  • webcam drain on network. proxy vs mirror?

Team reports

CH

  • Last week
    • AHS: submitted pump severity alert rationale. Received feedback from S. Ertel on what LBTI data needs to be collected by the AHS, and the high level conditional logic.
    • ARGOS: built arbitrator, some tweaks needed to get simulator running.
    • DMS: tested Monnit data ingest script for TLS 1.2 compatibility, ahead of iMonnit.com's planned blocking of < TLS 1.2.
  • This week
    • AHS: prototype the LBTI INDI driver.
    • ARGOS: tweaks needed to get simulator running.
    • DMS: get stub temperature page posting requests to its wsgi module.

PK

  • more IT7950 - see the IT, still evolving
  • new GCS, OAC - gathering for AGW debugging
  • Ice & LBC with Yang
  • moxa/mxio with Xianyu

DE

  • Polymer intro for shutdown tool
  • Made mods to Tyler's Vault user; no feedback yet
  • 'tis the season for password resets
  • Working with Igor re: hardware update on Luci computers
    • Plenty of room for another card
    • Aiming for SSD 2020
  • New storage volume on the summit ready to go
    • Xianyu, let me know what you need moved and how and I'll start writing a script

XZ

Last week:
  • Power loss reaction discussion and document with Guido
  • Fills AOCO WBS dictionary
  • Re-program MOXA (python data fromat transfer with Igor, C++ MXIO lib with Petr, finally, the python http was taken and tested in the lab)
This week (Monday and Tuesday):
  • Power backplane dicussion and review with Guido
  • AOCO review with Laura
Off-office from Dec 18 to Jan 12 (Vacation, holiday, and university closure)

YZ

  • Updated LBC sky flats Python program for Olga.
  • After multiple tests and dicussion with Petr, finally was able to compile TCSSendWavefront.c on Fedora30 with Ice3.7, GCC9, and current TCS Ice interface. Made a Github repo for CI build (to be pull requested).
  • Also worked on the Raritan Java client installation (made a Github repo to be pull requested).
  • Talked with Al about some new OSCO data collected and his newer C-based Strehl code (not yet working). Worked on testing several of the newer OSCO images with Steve Ertel's Python code with new Strehl results (to be analyzed by Al and Steve).
  • Will work on updating the OSCO people with a list of current available OSCO data.
  • Scheduled with Olga to discuss LBCFPIA code as well as trying to test the Raritan Java client settings.
  • Safety training.

MB

  • Last week
    • OC, WPC, Management
    • job post applicants review
    • DMS
      • url shortening without hash, thoughts
      • AdSec visuals toy experiments
    • SciArc mining, menu driven searches
    • AI/neural network donuts talk, no donuts, but brought back some ideas, intern suggestion from Al
  • This week
    • job post applicants review
    • AO data moving follow up
    • AHS deployment on hydrogen, update wiki doc, handover steps
    • DMS, neural networks
    • obsws, move fwd with rm507s
    • nsigui follow-up
    • AI follow up

SH

  • web mountain virtualization
  • webcams page -- everything broke on file locking, NFS is being weird in Tucson, modified to use less file locking, at jobs spun out, and some more modifications.
  • 507-5 display card replacement
  • Old web machines become PEPSI acquisition machines.

next:

  • p2v tucsondc1
  • new vm for web tucson... remove NFS dependency
  • move machine from AWS

IS

Last week:
  • LUCI
    • tail chasing ... it turns out that after 6 years of operations we never had a "this mask doesn't exist" message, when someone put a wrong mask on a script??
    • Apply amplitude magnification to cubes for LUCI2. The original code I put there didn't work for cubes. Did an emergency fix and deployed.
  • OT & PIT
    • Did some work assessing ETAs and in particular looking at the fits file loading in the OT target and pointing editor TPE
This week
  • Safety training
  • PIT
    • ongoing work on resubmission logic
  • OT
    • Fix LBC standards and MODS bias scripts with dgo
  • LUCI
    • Issue loading mask, wrong information in header files. Need to figure out why is that.
  • Set up my new laptop
Topic revision: r15 - 16 Dec 2019, MatthieuBec
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback